OpenStudy (camerondoherty):

The price of products may increase due to inflation and decrease due to depreciation. Derek is studying the change in the price of two products, A and B, over time. The price f(x), in dollars, of product A after x years is represented by the function below. f(x) = 10250(0.63)x Part A: Is the price of product A increasing or decreasing and by what percentage per year? Justify your answer. (5 points)

OpenStudy (camerondoherty):

I think its decreasing but I don't know by how much...

OpenStudy (anonymous):

is it (0.63)^x?

OpenStudy (anonymous):

@zscdragon is god

OpenStudy (camerondoherty):

and yes @Arnab09

OpenStudy (anonymous):

so it is decreasing as it is getting multiplied by 0.63 repeatedly, which is less than 1

OpenStudy (camerondoherty):

Thank you @Arnab09 and cn you help me ith part 2 of this question?

OpenStudy (camerondoherty):

and by how much is it decreasing exactly?

OpenStudy (anonymous):

if you are multiplying repeatedly by 0.63, how much is it decreasing per year? say the amount is y after one year it will be 0.63y so decrease = y-0.63y=0.37y so decrease per year= 0.37*100%=37%

OpenStudy (camerondoherty):

Thank You So Much! @Arnab09